### v2

# Real convergence and periodicity of pp–adic continued fractions

Abstract.
Continued fractions have been generalized over the field of pp–adic numbers, where it is still not known an analogue of the famous Lagrange’s Theorem. In general, the periodicity of pp–adic continued fractions is well studied and addressed as a hard problem. In this paper, we show a strong connection between periodic pp–adic continued fractions and the convergence to real quadratic irrationals. In particular, in the first part we prove that the convergence in $\mathbb{R}$ is a necessary condition for the periodicity of the continued fraction of a quadratic irrational in $\mathbb{Q}_{p}$. Moreover, we leave several conjectures on the converse, supported by experimental computations. In the second part of the paper, we exploit these results to develop a probabilistic argument for the non-periodicity of Browkin’s $p$-adic continued fractions. The probabilistic results are conditioned under the assumption of uniform distribution of the pp–adic digits of a quadratic irrational, that holds for almost all pp–adic numbers.
Key words and phrases:
Continued fractions; $p$-adic numbers, convergence, periodicity
2010 Mathematics Subject Classification:
11J70; 11D88; 11Y65; 12J25
Department of Mathematical Sciences Giuseppe Luigi Lagrange,
Politecnico di Torino, Corso Duca degli Abruzzi 24, Torino
giuliano.romeo@polito.it
1. Introduction
Continued fractions are expressions of the form
$a_{0}+\cfrac{1}{a_{1}+\cfrac{1}{a_{2}+\ddots}},$ (1)
and they are a very classical and powerful tool in number theory. They have been studied and used throughout the centuries in all areas of mathematics, mainly due to their excellent properties of approximation. Continued fractions have been employed, for example, to prove the irrationality of $\pi$ [21], to construct transcendental numbers [1, 25], to find indecomposable algebraic integers in quadratic number fields [16] and for the factorization of integers [24]. To have a fairly complete introduction on the classical theory of continued fractions, see [19, 34, 40]. In the field of real numbers, continued fractions completely characterize rational numbers and quadratic irrationals. A real number $\alpha\in\mathbb{R}$ is rational if and only if its continued fraction is finite. The latter is an easy consequence of the finiteness of the Euclidean division algorithm in $\mathbb{Z}$. Moreover, $\alpha\in\mathbb{R}$ is a quadratic irrational, i.e. a root of an irreducible rational polynomial of degree $2$, if and only if its continued fraction is eventually periodic. The latter is the famous Lagrange’s Theorem [20]. Because of their optimal properties, continued fractions are highly studied in mathematics and there exist several generalizations. The topic of this paper is the study of continued fractions in the field of pp–adic numbers $\mathbb{Q}_{p}$, introduced by Mahler in [26]. See [36] for a recent survey on the theory of pp–adic continued fractions. The problem of defining a continued fraction algorithm in $\mathbb{Q}_{p}$ sharing all the optimal properties enjoyed by classical continued fractions is still open. In particular, it is not known an algorithm that provides an eventually periodic continued fraction for every quadratic irrational in $\mathbb{Q}_{p}$. Hence, pp–adic continued fractions miss, up to now, an analogue of Lagrange’s Theorem for classical continued fractions.
The first definitions date back around the 1970s by Ruban [37], Schneider [38] and Browkin [8]. The reason of the non-existence of a unique standard algorithm is that there is not a unique choice for the floor function of a pp–adic number. The choices of Ruban, Schneider and Browkin are all meaningful with respect to different points of view. However, Ruban’s and Schneider’s continued fractions are not finite for every rational number, but they can also be infinite periodic [10, 23]. Moreover, these two algorithms do not provide a periodic continued fraction for every pp–adic quadratic irrational. In fact, the analogues of Lagrange’s Theorem have been disproved in [39] for Schneider’s algorithm and in [14] for Ruban’s algorithm. In 1978 and 2000, Browkin [8, 9] came out with the definition of two pp–adic continued fractions algorithms (we call them Browkin I and Browkin II) that terminate if and only if the input is a rational number. The proof of finiteness for Browkin I is contained in Browkin’s first paper [8], while for Browkin II was conjectured by Browkin in [9] and proved later on by Barbero, Cerruti and Murru in [2]. Browkin’s first algorithm is really similar to Ruban’s one, but it allows to have also negative numbers as partial quotients. In fact, Browkin’s choice is to take, in the pp–adic expansion of an element of $\mathbb{Q}_{p}$, the representatives modulo $p$ in $\{-\frac{p-1}{2},\ldots,\frac{p-1}{2}\}$ instead of $\{0,\ldots,p-1\}$. We will refer to these kind of algorithms, choosing representatives in the symmetric interval $\{-\frac{p-1}{2},\ldots,\frac{p-1}{2}\}$, as Browkin-type algorithms.
The problem of understanding if Browkin-type continued fractions are always periodic for pp–adic quadratic irrationals or not is still open, and it is usually addressed as a hard problem. The main difference with Ruban’s and Schneider’s cases is the presence of negative partial quotients in the continued fraction expansion (see Section 3 for more details). The periodicity of Browkin I has been studied by Bedocchi in [4, 5] and by Capuano, Murru and Terracini in [12]. The periodicity of Browkin II has been deepened in [2, 29]. In particular, Browkin II computationally appears to outperform Browkin I in the number of periodic continued fractions provided for quadratic irrational numbers. Moreover, the convergence conditions proved in [30] allowed to define a new algorithm in [28], that is introduced in Section 2 as Algorithm (12). The algorithm defined in [28] is similar to Browkin II but it improves its periodicity properties from both theoretical and computational points of view. Other algorithms have been defined and studied in [3, 15] for standard pp–adic continued fractions and in [31, 32, 33] for the multidimensional case. See also [11, 13] for other generalizations of pp–adic continued fractions. All these algorithms improve and generalize the known algorithms for pp–adic continued fractions from several different points of view.
The main object of study of the present paper is the periodicity of the Browkin-type algorithms defined in [8, 9, 28] and introduced in the next section, i.e. Browkin I, Browkin II and Algorithm (12). Although Lagrange’s Theorem has not been proved or disproved for any these three algorithms, it is widely believed to fail (see, for example, the computational analysis performed in [28]). However, as we discuss in Section 3, it does not exist at the present time an argument to deal efficiently with the periodicity of Browkin-type continued fractions. The main reason is that the argument carried out in the pioneering work of Capuano, Veneziano and Zannier [14] for the periodicity of Ruban’s continued fractions, is specific for continued fractions with positive partial quotients. For these reasons, at the present time, no quadratic irrational has been proved to have non-periodic pp–adic continued fraction with a Browkin-type algorithm.
In this paper, we explore a different and novel argument for the non-periodicity of Browkin-type continued fractions. In particular, we highlight a strong connection between the periodicity of a pp–adic continued fraction and the convergence in $\mathbb{R}$ to a real quadratic irrational. The core result and the starting point for our analysis is the following necessary condition for the periodicity.
Proposition 1.
Let $\alpha$ be a quadratic irrational that has both a real image $\alpha^{(r)}$ and a pp–adic image $\alpha^{(p)}$. If the pp–adic continued fraction $a_{0},a_{1},\ldots]$ of $\alpha^{(p)}$ obtained with Browkin I, Browkin II or Algorithm (12) is eventually periodic, then $a_{0},a_{1},\ldots]$ converges in $\mathbb{R}$ to either $\alpha^{(r)}$ or $\overline{\alpha}^{(r)}$ .
Proposition 1 can be proved by using a result for the convergence of generalized continued fractions with complex partial quotients [40, Theorem 8.1], based on the results of Lane [22], and it is discussed more thoroughly in Section 4. The idea is that, by requiring the periodicity, the continued fraction is a root of a polynomial of degree $2$ over $\mathbb{Q}$, and this is true both embedding the continued fraction in $\mathbb{Q}_{p}$ and in $\mathbb{R}$. Moreover, Proposition 1 was also conjectured in a short note by Moore, that is not anymore available online.
In Section 3, we highlight as working by embedding pp–adic continued fractions into the field of real numbers has been, up to now, one of the most useful techniques to obtain effective results. In fact, it is the core idea of the nice characterization of the periodicity of Ruban’s continued fractions proved in [14]. Also the strategy for the proof of finiteness of Browkin I for all rational numbers, carried out in [8], exploits an inequality for the Euclidean absolute value of the partial quotients. Moreover, it is also the underlying idea of the algorithms defined in [15], producing periodic pp–adic continued fractions of quadratic irrationals for small values of $p$.
From the experimental computations, it seems that the convergence in $\mathbb{R}$ to either $\alpha^{(r)}$ or $\overline{\alpha}^{(r)}$ of the pp–adic continued fraction of $\alpha^{(p)}$ is also a sufficient condition for periodicity, i.e. also the converse of Proposition 1 is true. In fact, in all the cases where the continued fraction $a_{0},a_{1},\ldots]$ of a quadratic irrational $\alpha^{(p)}$ seems to be non-periodic, as a period is not detected for many steps, then $a_{0},a_{1},\ldots]$ seems converging in $\mathbb{R}$ very neatly to a limit different from $\alpha^{(r)}$ (see the computational analysis performed in Section 6). For this reason, we leave the following conjecture.
Conjecture 2.
Let $a_{0},a_{1},\ldots]$ be the Browkin I, Browkin II or Algorithm (12) continued fraction of a $p$-adic quadratic irrational $\alpha^{(p)}$ that can be embedded in the real numbers. Let us call $\alpha^{(r)}$ the image of $\alpha^{(p)}$ in $\mathbb{R}$. Then, the continued fraction $a_{0},a_{1},\ldots]$ is eventually periodic if and only if it converges with respect to the Euclidean absolute value to either $\alpha^{(r)}$ or $\overline{\alpha}^{(r)}$.
In the next sections, we provide several examples in support of Conjecture 2. After many steps, the continued fraction of a quadratic irrational $\alpha^{(p)}$ tends to oscillate less and less around a real limit, and the latter seems to be, in the cases where no periodicity is observed, really far from $\alpha^{(r)}$ and $\overline{\alpha}^{(r)}$. Moreover, a big deviation between consecutive convergents must correspond to small values of the partial quotients $a_{n}$. Small values of the partial quotients are very unlikely, especially for large values of the prime $p$. In fact, it is widely believed that the digits of the pp–adic expansions of irrational numbers are uniformly distributed in $\mathbb{Z}/p\mathbb{Z}$. Therefore, in Section 5, we develop a probabilistic argument for the non-periodicity of pp–adic continued fractions obtained with Browkin I, and similar arguments can be applied for Browkin II and Algorithm (12). The probabilistic results are conditioned under the following assumption.
Assumption 3.
Let us consider $\alpha=\sum\limits_{n=r}^{+\infty}c_{n}p^{n}$ be a random pp–adic quadratic irrational. Then, for all $n\geq r$ and for all $k\in\{0,\ldots,p-1\}$,
$\mathbb{P}(c_{n}=k)=\frac{1}{p},$
with respect to the Haar measure. In other words, the coefficients of the $p$-adic expansion of $\alpha$ are uniformly distributed in the set $\{0,\ldots,p-1\}$.
The reasonability of Assumption 3 is based on Borel’s normal number theorem [6]. In 1909, Borel proved that almost all real numbers are normal, i.e. the set of non-normal numbers has Lebesgue measure 0 in $\mathbb{R}$. A real number $\alpha$ is (absolutely) normal if, for any integer $b\geq 2$, every digit of its expansion in base $b$ has density $\frac{1}{b}$. It is known that Borel’s argument, exploiting Borel-Cantelli Lemma, can be applied equivalently to the pp–adic expansion of pp–adic numbers. Therefore, almost all pp–adic numbers are normal. In particular, all irrational numbers, both in $\mathbb{R}$ and in $\mathbb{Q}_{p}$, are conjectured to be normal. However, apart from Borel’s existence result, only few numbers have been proved to be normal. For example, there is not a proof that $\sqrt{2}$, $\pi$ or $e$ are normal, although it is strongly considered true.
In Theorem 26 of Section 5, under Assumption 3, we prove that the expected value of any Browkin I partial quotient $a$ is
$\mathbb{E}(|a|)=\frac{p}{4}\left(1-\frac{1}{p^{2}(p^{2}+p+1)}\right),$ (2)
therefore it grows linearly with the prime $p$. This means that the convergents are really unlikely to deviate after several steps, especially for large $p$. For instance, see Example 28 at the end of Section 5, where the 78237823–adic continued fraction of $\sqrt{15648}$ has the convergent
$\frac{A_{10000}}{B_{10000}}\approx 3337.$ (3)
If its Browkin I expansion became periodic, then, by Proposition 1, it should be convergent to $\sqrt{15648}\approx 125$. However, it is really unlikely for the subsequent convergents to deviate from (3), as $\mathbb{E}(|a_{n}|)\approx 1955.75$.
Notice that the growth of the denominators $|B_{n}|$ tends to be exponential, as the main term is $|a_{0}a_{1}\ldots a_{n}|$. Therefore, we leave the following conjecture.
Conjecture 4.
Let $\{|B_{n}|\}_{n\in\mathbb{N}}$ be the sequence of denominators of convergents for Browkin I, Browkin II and Algorithm (12) $p$-adic continued fractions, where $p\geq 5$. Then $|B_{n}|$ tends to $\infty$ with probability $1$.
2. Preliminaries and main algorithms
In this section we recall some basic facts from the theory of continued fractions and pp–adic numbers. For more results and background we refer the reader to [34, 40] for continued fractions and to [18] for pp–adic numbers.
In the following, let $p$ be an odd prime number. We denote by $v_{p}(\cdot)$ the pp–adic valuation and by $|\cdot|$ and $|\cdot|_{p}$, respectively, the Euclidean absolute value and the pp–adic absolute value. We denote the continued fraction expansion
$a_{0}+\cfrac{1}{a_{1}+\ddots}$ (4)
by $a_{0},a_{1},\ldots]$. The coefficients $a_{n}$ are called partial quotients and they are, for our purposes, rational numbers. For all $n\in\mathbb{N}$, the rational number
$\frac{A_{n}}{B_{n}}=[a_{0},a_{1},\ldots,a_{n}]=a_{0}+\cfrac{1}{a_{1}+\ddots+\cfrac{1}{a_{n-1}+\cfrac{1}{a_{n}}}},$ (5)
corresponding to the continued fraction stopped at the $n$-th term, is called $n$-th convergent of the continued fraction. It is not hard to prove that the sequences $\{A_{n}\}_{n\in\mathbb{N}}$ and $\{B_{n}\}_{n\in\mathbb{N}}$ in (5) satisfy the following recursions:
$\begin{cases}A_{0}=a_{0},\\ A_{1}=a_{1}a_{0}+1,\\ A_{n}=a_{n}A_{n-1}+A_{n-2},\ \ n\geq 2,\end{cases}\begin{cases}B_{0}=1,\\ B_{1}=a_{1},\\ B_{n}=a_{n}B_{n-1}+B_{n-2},\ \ n\geq 2.\end{cases}$ (6)
Moreover, they satisfy, for all $n\in\mathbb{N}$, the relation:
$A_{n}B_{n+1}-A_{n+1}B_{n}=(-1)^{n+1}$ (7)
The standard algorithm to express a real number $\alpha$ through a simple continued fraction $a_{0},a_{1},\ldots]$ works as follows, starting with $\alpha_{0}=\alpha$:
$\begin{cases}a_{n}=\lfloor\alpha_{n}\rfloor\\ \alpha_{n+1}=\frac{1}{\alpha_{n}-a_{n}},\end{cases}$ (8)
where $\lfloor\alpha_{n}\rfloor$ denotes the integer part of $\alpha$. The elements $\alpha_{n}$ are called complete quotients. If $\alpha_{n}=a_{n}$ for some $n\in\mathbb{N}$, then the algorithm terminates and the continued fraction is finite.
One of the main strategies for introducing continued fractions in $\mathbb{Q}_{p}$ has been to search for an analogue of the floor function $\lfloor\cdot\rfloor$ that is meaningful inside $\mathbb{Q}_{p}$, in order to emulate the standard algorithm (8). The first algorithms have been proposed by Ruban [37] and Schneider [38]. Browkin’s first algorithm [8] is very similar to Ruban’s algorithm, with the exception that the representatives of $\mathbb{Z}/p\mathbb{Z}$ are chosen in $\{-\frac{p-1}{2},\ldots,\frac{p-1}{2}\}$ instead of $\{0,\ldots,p-1\}$. This small variation is fundamental because Browkin’s algorithm produces a finite continued fraction for each rational number, while Ruban’s algorithm can also be periodic over the rationals. Given $\alpha=\sum\limits_{i=-r}^{+\infty}c_{i}p^{i}\in\mathbb{Q}_{p}$, with $c_{i}\in\{-\frac{p-1}{2},\ldots,\frac{p-1}{2}\}$, Browkin defines the floor function $s:\mathbb{Q}_{p}\rightarrow\mathbb{Q}$ as
$s(\alpha)=\sum\limits_{i=-r}^{0}c_{i}p^{i},$ (9)
and $s(\alpha)=0$ if $r<0$. Given the floor function $s$, the algorithm is the same as (8). Browkin’s first algorithm, that we call Browkin I, work as follows. At the first step $\alpha_{0}=\alpha$ and, for all $n\geq 0$,
$\begin{cases}a_{n}=s(\alpha_{n})\\ \alpha_{n+1}=\frac{1}{\alpha_{n}-a_{n}}.\end{cases}$ (10)
If at some point $\alpha_{n}=a_{n}$, then the algorithm stops and $\alpha=[a_{0},\ldots,a_{n}]$, i.e. $\alpha$ has finite Browkin I continued fraction.
In [9], Browkin defined another floor function that is similar to the first function $s$, but excluding the summand with valuation zero. For $\alpha=\sum\limits_{i=-r}^{+\infty}c_{i}p^{i}\in\mathbb{Q}_{p}$, with $c_{i}\in\{-\frac{p-1}{2},\ldots,\frac{p-1}{2}\}$, the second floor function is the function $t:\mathbb{Q}_{p}\rightarrow\mathbb{Q}$, such that
$t(\alpha)=\sum\limits_{i=-r}^{-1}c_{i}p^{i},$
and $t(\alpha)=0$ if $r\leq 0$. The second algorithm, that we call Browkin II, works on an input $\alpha$ as follows. At the first step $\alpha_{0}=\alpha$ and, for all $n\geq 0$,
$\begin{cases}a_{n}=s(\alpha_{n})\ \ \ \ \ &\textup{if}\ n\ \textup{even}\\ a_{n}=t(\alpha_{n})&\textup{if}\ n\ \textup{odd}\ \textup{and}\ v_{p}(\alpha_{n}-t(\alpha_{n}))=0\\ a_{n}=t(\alpha_{n})-sign(t(\alpha_{n}))&\textup{if}\ n\ \textup{odd}\ \textup{and}\ v_{p}(\alpha_{n}-t(\alpha_{n}))\neq 0\\ \alpha_{n+1}=\frac{1}{\alpha_{n}-a_{n}}.\end{cases}$ (11)
If at some point $\alpha_{n}=a_{n}$, then the algorithm stops and $\alpha=[a_{0},\ldots,a_{n}]$, i.e. $\alpha$ has finite Browkin II continued fraction. One consequence of the alternation of the two functions is that, in Browkin II, all the even partial quotients are by constructions integers and all the odd partial quotients are non-integer rationals. In [28], a modification of Browkin II has been defined, that does not use the sign function. For all $\alpha_{0}\in\mathbb{Q}_{p}$, the algorithm works as follows:
$\begin{cases}a_{n}=s(\alpha_{n})\ \ \ \ \ &\textup{if}\ n\ \textup{even}\\ a_{n}=t(\alpha_{n})&\textup{if}\ n\ \textup{odd}\\ \alpha_{n+1}=\frac{1}{\alpha_{n}-a_{n}}.\end{cases}$ (12)
It has been proved in [28] that Algorithm (12) improves the periodicity properties of Browkin’s second algorithm. For more details and a computational analysis on the performance of Browkin I, Browkin II and Algorithm (12), see [28].
3. General remarks on the periodicity
Lagrange’s Theorem states that the standard continued fraction expansion of a real quadratic irrational number becomes eventually periodic. The most natural technique is to divide the proof in two steps. First, a theorem of Galois [17] characterizes the purely periodic continued fractions.
Theorem 5 (Galois’ Theorem).
The continued fraction expansion of a quadratic irrational $\alpha\in\mathbb{R}$ is purely periodic if and only if it is reduced, that is $\alpha>1$ and $1<\overline{\alpha}<0$, where $\overline{\alpha}$ is the conjugate of $\alpha$ over $\mathbb{Q}$.
The idea to prove Galois’ Theorem is that if a complete quotient
$\alpha_{n}=\frac{P_{n}+\sqrt{D}}{Q_{n}}$ (13)
is reduced, also $\alpha_{n+1}$ is reduced. Then, it is not hard to see that requiring $|\alpha_{n}|>1$ and $1<\overline{\alpha}_{n}<0$ gives rise to only finitely many different choices for $P_{n}$ and $Q_{n}$. Therefore, a repetition $\alpha_{n}=\alpha_{n+k}$, $k\geq 1$, is met at some point, and the continued fraction is periodic. The standard proof of Lagrange’s Theorem consists in proving that the expansion of every quadratic irrational eventually meets a reduced quadratic irrational among its complete quotients. Hence, it starts to be periodic at some point.
In $\mathbb{Q}_{p}$ there are some characterizations of purely periodic continued fractions, that are really similar to Galois’ Theorem both in the statement and in the proof.
Theorem 6 ([14],[4]).
Let $\alpha\in\mathbb{Q}_{p}$ having a periodic Ruban’s (or Browkin’s) continued fraction. Then the continued fraction is purely periodic if and only if $|\alpha|_{p}>1$ and $|\overline{\alpha}|_{p}<1$.
Theorem 7 ([28]).
Let $\alpha\in\mathbb{Q}_{p}$ having a periodic continued fraction with Algorithm (12). Then the continued fraction is purely periodic if and only if $|\alpha|_{p}\geq 1$ and $|\overline{\alpha}|_{p}<1$.
In Theorems 6 and 7, the hypothesis of starting from a periodic continued fraction can not be removed. This is in great contrast with Galois’ Theorem, where the conditions on the absolute values of $\alpha$ and $\overline{\alpha}$ are also sufficient for the periodicity.
The main problem here is the use of the pp–adic absolute value instead of the Euclidean one. In fact, unlike the real case, there are infinitely many pp–adic quadratic irrationals $\alpha$ satisfying $|\alpha|_{p}>1$ (or $\geq 1$) and $|\overline{\alpha}|_{p}<1$. Therefore, being reduced is not a sufficient condition for periodicity in $\mathbb{Q}_{p}$.
In [14], Capuano, Veneziano and Zannier proved that there are only finitely many quadratic irrational with same discriminant $D$ having a periodic Ruban’s continued fraction. This allowed to prove an effective criterion for determining in a finite number of steps if a given quadratic irrational becomes periodic or not.
The idea carried out in [14] exploits a “real” argument more than a “pp–adic” argument. In fact, it is used the real image of the quadratic irrational $\alpha$ in $\mathbb{R}$. If the pp–adic quadratic irrational number
$\alpha=\frac{P+\sqrt{D}}{Q}$
has got a purely periodic expansion with Ruban’s algorithm, then $\alpha=\alpha_{k}$ for some $k$, hence
$\alpha=\frac{\alpha_{k}A_{k-1}+A_{k-2}}{\alpha_{k}B_{k-1}+B_{k-2}}=\frac{\alpha A_{k-1}+A_{k-2}}{\alpha B_{k-1}+B_{k-2}},$
and it is a root of the polynomial
$B_{k-1}\alpha^{2}-(A_{k-1}-B_{k-2})\alpha-A_{k-2}=0.$
Since all the partial quotients in Ruban’s expansion are positive, then the numerators $A_{n}$ and the denominators $B_{n}$ are positive for all $n\in\mathbb{N}$. Therefore, if $\alpha$ has a periodic expansion, then
$\alpha\overline{\alpha}=-\frac{A_{k-2}}{B_{k-1}}<0.$ (14)
If $\alpha$ is purely periodic, then all its complete quotients $\alpha_{n}$ are purely periodic and hence satisfy (14). However, this means that
$\alpha_{n}\overline{\alpha}_{n}=\frac{P_{n}^{2}-D}{Q_{n}^{2}}<0,$ (15)
for all $n\in\mathbb{N}$. It is not hard to see that (15) gives an upper bound on both $P_{n}$ and $Q_{n}$, depending only on $D$, similar to those obtained in the real case. Therefore, there are only finitely many different complete quotients $\alpha_{n}$ for a pp–adic number $\alpha$ having a purely periodic Ruban’s expansion. This gives an effective practical criterion for determining the periodicity or non-periodicity of Ruban’s continued fractions, since in finite time (practically determinable) there is either a repetition or a complete quotient not satisfying (14).
Unfortunately, a similar technique is not applicable for Browkin-type algorithms. In fact, since Browkin’s partial quotients can also be negative, then (14) is not a necessary condition for the periodicity. This is one of the main reasons why the problem of determining if an analogue of Lagrange’s Theorem holds or fails is still unsolved. The best result in this sense is the following, and it has been proved by Capuano, Murru and Terracini in [12] for Browkin I.
Proposition 8 ([12], Proposition 4.8).
Let $\alpha\in\mathbb{Q}_{p}$ be a quadratic irrational and for every $n>0$ denote by $\xi_{n}$ and $\xi^{\prime}_{n}$ the two images of $\alpha_{n}$ in $\mathbb{C}$ and $t=\lfloor\sqrt{\Delta}\rfloor$. Assume that $\exists n_{0}>0$ such that $\xi_{n}\xi^{\prime}_{n}<0$ for every $n\in\left[n_{0},n_{0}+K\right]$, where $K=K(t)$ is a constant depending only on $t$. Then the Browkin I continued fraction of $\alpha$ is periodic with a period of length at most $K$.
The proof of Proposition 8 exploits an argument really similar to the one carried out in [14], starting from Equation (14).
Remark 9.
In order to apply the argument of [14] to obtain a criterion for the periodicity of Browkin-type pp–adic continued fractions, we do not necessarily need (14). It would suffice to have, for some $K>0$,
$\frac{A_{k-1}}{B_{k}}>-K,$ (16)
so that
$\alpha\overline{\alpha}=-\frac{A_{k-1}}{B_{k}}<K.$ (17)
This would lead to some bound on the sequences $P_{n}$ and $Q_{n}$ and hence to an effective criterion for the periodicity of Browkin’s continued fractions. However, even if the constant $K$ can be arbitrarily large, it is not easy (or probably not possible at all) to find a $p$-adic quadratic irrational $\alpha$ such that its convergents satisfy (16).
For all these reasons, there is not an “easy” way to prove non-periodicity for Browkin I up to now, and similar reasonings can be performed for Browkin II or Algorithm (12). At the present time, no quadratic irrational has been proved to have non-periodic Browkin I continued fraction, although it is largely believed to fail Lagrange’s Theorem (see, for example, the experimental computations in [28]).
Before delving into the real convergence of pp–adic continued fractions, let us express few other words on the effectiveness of the real approach for pp–adic continued fractions. Browkin’s proof [8] that every rational numbers has a finite Browkin I continued fraction could not be carried out without an argument controlling the Euclidean size of the partial quotients inside $\mathbb{R}$. In fact, the standard proof of finiteness for rational numbers (that also inspired all the subsequent proofs of finiteness in [2, 3, 15, 28, 30]) consists in producing a strictly decreasing sequence of integers in Euclidean absolute value. It is not possible to use a similar argument inside $\mathbb{Q}_{p}$, since there are infinite sequences of integers that are strictly decreasing in pp–adic absolute value. Hence, all the proofs of finiteness for the pp–adic continued fractions of rational numbers that are known so far, are performed by embedding the continued fractions inside $\mathbb{R}$. It is remarkable also the technique carried out in [27] for the archimedean setting and then in [13] for the $p$-adic one, in order to prove finiteness for continued fractions defined over quadratic number fields. The arguments use Weil height and Northcott’s theorem. This is an interesting approach because Weil height takes into account all absolute values of an algebraic number, not just the Euclidean or the $p$-adic one for a fixed $p$.
In [15], Deng and Wang defined a continued fraction algorithm following an idea similar to Browkin’s proof of finiteness, but for proving the periodicity. They introduced two novel floor functions $\tilde{s}$ and $\tilde{t}$ that minimize the Euclidean absolute value of $\alpha-\tilde{s}(\alpha)$ and $\alpha-\tilde{t}(\alpha)$. This machinery allowed to find some nice inequalities for the numerators $P_{n}$ and the denominators $Q_{n}$ of the complete quotients $\alpha_{n}$. These inequalities led, for small values of the prime $p$, to a strictly decreasing sequence involving $P_{n}$ and $Q_{n}$. Therefore, for these small values of $p$, the pp–adic continued fraction of every quadratic irrational is eventually periodic, using their algorithm. This is the first proof of periodicity of an algorithm in some $\mathbb{Q}_{p}$ for every pp–adic quadratic irrational (at least for those having positive discriminant, hence a real embedding). A periodic continued fraction expression for every quadratic irrational $\alpha$ having both a pp–adic and a real interpretation was already defined in [2]. In [2], the authors exploited the simultaneous convergence to $\alpha$ in $\mathbb{R}$ and $\mathbb{Q}_{p}$ of the Rédei rational functions $R_{n}$ [35] and provided a periodic continued fraction having exactly the $R_{n}$ as convergents.
However, the important weakness of these algorithms is that they work only for quadratic irrationals, and in particular for those that can be embedded in both $\mathbb{Q}_{p}$ and $\mathbb{R}$. Moreover, they require the explicit knowledge of the minimal polynomial of the quadratic irrational. This is in great contrast with the spirit of Lagrange’s Theorem for classical continued fractions. In fact, the standard algorithm works on a generic $\alpha\in\mathbb{R}$ and it is eventually periodic if and only if $\alpha$ is a quadratic irrational.
4. Real convergence of pp–adic continued fractions
The problem in full generality could be stated as: when does a continued fraction $a_{0},a_{1},\ldots]$ converge both in the pp–adic and in the real topology? Not always, of course. For example, the well known continued fraction
$\overline{1}]=\frac{1+\sqrt{5}}{2},$ (18)
converges in $\mathbb{R}$ to the golden mean, but does not converge in $\mathbb{Q}_{p}$. Moreover, also a continued fraction $a_{0},a_{1},\ldots]$ with rational partial quotients that converges in $\mathbb{Q}_{p}$ not necessarily converges in $\mathbb{R}$ (not even in $\mathbb{C}$). For example
$\left[\frac{1}{p},\frac{1}{p^{2}},\frac{1}{p^{3}},\ldots\right]$
converges to a pp–adic number since $v_{p}(a_{n})<0$ for all $n$, but the series
$\sum\limits_{n=0}^{\infty}|a_{n}|=\sum\limits_{n=0}^{\infty}\frac{1}{p^{n+1}}=\frac{1}{p-1},$
converges in Euclidean absolute value. Therefore, the continued fraction $a_{0},a_{1},\ldots]$ diverges to $\infty$ in the Euclidean setting, due to a known result that we recall here.
Proposition 10 ([40], Theorem 6.1).
If the real continued fraction $a_{0},a_{1},\ldots]$ converges to a real number, then the series $\sum\limits_{n=0}^{+\infty}|a_{n}|$ diverges to $\infty$.
Convergence is usually easier for periodic continued fractions and there are much more precise results than the non-periodic case. However, as we have seen in (18), periodic continued fractions that converge in $\mathbb{R}$ not necessarily converge in $\mathbb{Q}_{p}$. Also periodic pp–adic continued fractions can be not convergent with respect to the Euclidean absolute value. For example, if we consider $\overline{\frac{2}{5},-\frac{2}{5}}]$, then
$\alpha=\frac{\alpha_{2}A_{1}+A_{0}}{\alpha_{2}B_{1}+B_{0}}=\frac{\frac{21}{25}\alpha+\frac{2}{5}}{-\frac{2}{5}\alpha+1}.$ (19)
This means that $\alpha$ is a root of $5x^{2}-2x+5$, that has negative discriminant $\Delta=-96$. Therefore, it should converge to a non-real complex number, which is clearly not possible, as the convergents are rationals and $\mathbb{R}$ is a complete field.
In the following, we delve into the case of our interest, that is when the continued fractions are obtained starting from an element $\alpha$. In order to properly speak about real and pp–adic convergence to $\alpha$ of a continued fraction, $\alpha$ must be embedded in both the reals and the pp–adics. Therefore we introduce some context and notation for the important cases of our study, the quadratic irrationals.
A quadratic irrational is a root of an irreducible polynomial
$f(x)=ax^{2}+bx+c\in\mathbb{Q}[x]$
of degree $2$ with rational coefficients. In our analysis, we deal with polynomials having a root both in $\mathbb{R}$ and in $\mathbb{Q}_{p}$. This means that the discriminant $D=b^{2}-4ac$ is positive, so that the equation $f(x)=0$ is solvable in $\mathbb{R}$, and it is a quadratic residue modulo $p$, so that the equation $f(x)=0$ is solvable in $\mathbb{Q}_{p}$. We call $\alpha^{(r)}$ and $\overline{\alpha}^{(r)}$ its roots in $\mathbb{R}$ and we call $\alpha^{(p)}$ and $\overline{\alpha}^{(p)}$ its roots in $\mathbb{Q}_{p}$.
Capuano, Veneziano and Zannier [14] proved the following result for Ruban’s continued fractions.
Theorem 11 ([14], Proposition 4.5).
The Ruban’s continued fraction of a pp–adic quadratic irrational $\alpha^{(p)}$ always converges in $\mathbb{R}$ to some real number $\xi$.
Ruban’s continued fractions are not always periodic for quadratic irrationals. The periodic expansion of a pp–adic quadratic irrational $\alpha^{(p)}$, with a technique similar to Proposition 1, can be showed to converge in $\mathbb{R}$ to either $\alpha^{(r)}$ or $\overline{\alpha}^{(r)}$. However, when Ruban’s continued fraction of $\alpha^{(p)}$ is non-periodic, we know that it converges in $\mathbb{R}$, but we do not know to which real number. The authors of [14] left the following conjecture.
Conjecture 12 ([14], Section 4.2).
Ruban’s continued fraction of a quadratic irrational is either periodic or it converges in $\mathbb{R}$ to a transcendental number.
This conjecture is quite reasonable, since algebraic numbers are countable and hence have measure 0 in $\mathbb{R}$. Therefore, we expect a continued fraction having an “unknown” limit to converge to a transcendental number.
For Browkin’s continued fractions, we have experimentally observed similar results, for quadratic irrationals having a real image (see Section 6). In the following we speak of either Browkin I or Browkin II or Algorithm (12), since they show similar behaviours of real convergence. We leave the following conjecture, that is similar to Theorem 11 for Browkin-type continued fractions.
Conjecture 13.
Browkin I, Browkin II or Algorithm (12) $p$-adic continued fractions of quadratic irrationals that can embedded into real numbers always converge in $\mathbb{R}$.
In fact, it computationally appears that Browkin’s and Algorithm (12) continued fractions converge, both in the periodic and the non-periodic cases, to some real number (see computations in Section 6). The problem in generalizing the proof of Theorem 11 is again the presence of negative partial quotients.
Remark 14.
Notice the difference between the statement of Theorem 11 and our Conjecture 13. In fact, in Theorem 11 the real convergence is proved for all quadratic irrationalities, also in the case where they can not be embedded in the reals, i.e. they have negative discriminant. This implies, in particular, that imaginary quadratic irrationals can not have a periodic Ruban’s continued fraction. Instead, Conjecture 13 is stated only for quadratic irrationalities having positive discriminant. In fact, it is false for imaginary quadratic irrationals as, for example,
$\sqrt{-6}=\left[2,-\frac{7}{5},\overline{\frac{8}{5},-\frac{12}{5}}\right],$
has a periodic Browkin I continued fraction in $\mathbb{Q}_{5}$ and, hence, can not converge to a real number. On the contrary, Browkin I continued fractions of imaginary quadratic irrationals for which periodicity is not observed, seem to converge to a real number.
Now we prove Proposition 1, stating that, in the periodic case, the pp–adic continued fraction converges to the “same” quadratic irrational in $\mathbb{R}$. In fact, we show that periodic pp–adic continued fractions of quadratic irrationals that have embeddings both in $\mathbb{R}$ and in $\mathbb{Q}_{p}$, converge simultaneously in $\mathbb{R}$ and in $\mathbb{Q}_{p}$ to roots of the same minimal polynomial. Before proving Proposition 1, we recall the following classical and very general result for the convergence of periodic complex continued fractions in the Euclidean setting. The proof can be found in Wall’s book [40] or in Lane’s original paper [22].
Theorem 15 ([40], Theorem 8.1).
Let $\overline{a_{0},\ldots,a_{k-1}}]$ be a periodic continued fraction with complex partial quotients and let us call $\alpha_{1}$ and $\alpha_{2}$ the two roots of the polynomial relation of degree $2$ obtained by the periodicity. Then the continued fraction converges in $\mathbb{C}$ if and only if $\alpha_{1}$ and $\alpha_{2}$ are finite and one of the following conditions holds:
  1. i)
    $\alpha_{1}=\alpha_{2}$,
  2. ii)
    $\left|\alpha_{2}-\frac{A_{k-1}}{B_{k-1}}\right|>\left|\alpha_{1}-\frac{A_{k-1}}{B_{k-1}}\right|$, and $\frac{A_{n}}{B_{n}}\neq\alpha_{2}$ for $n=0,\ldots,k-2$.
We now prove Proposition 1 by verifying that Condition ii)ii) of Theorem 15 is satisfied by Browkin I, Browkin II and Algorithm (12) pp–adic continued fractions. Alternatively, it would be possible to use a technique similar to [7, Theorem 4.3].
Proof of Proposition 1.
Let us consider the periodic pp–adic continued fraction
$a_{0},\ldots,a_{h-1},\overline{a_{h},\ldots,a_{h+k-1}}],$
obtained for the quadratic irrational $\alpha^{(p)}$ by means of either Browkin I, Browkin II or Algorithm (12). We know that $a_{0},a_{1},\ldots]$ converges to $\alpha^{(p)}$ in $\mathbb{Q}_{p}$ by the correctness of the algorithms. Now we show that its periodic part satisfies the hypothesis ii)ii) of Theorem 15 and, hence, that $a_{0},a_{1},\ldots]$ converges to either $\alpha^{(r)}$ or $\overline{\alpha}^{(r)}$ in $\mathbb{R}$. In fact, the convergence is clearly not affected by the presence of the pre-period, and the periodicity gives raise to the same minimal polynomial of $\alpha^{(p)}$. Therefore, in the following we consider for simplicity $h=0$, i.e. the purely periodic continued fraction $\overline{a_{0},\ldots,a_{k-1}}]$. Notice that in order to satisfy condition ii)ii), since all the partial quotients are real numbers, it is sufficient to prove that it never happens that
$\alpha_{2}^{(r)}-\frac{A_{k-1}}{B_{k-1}}=-\alpha_{1}^{(r)}+\frac{A_{k-1}}{B_{k-1}},$ (20)
where $\alpha_{1}^{(r)},\alpha_{2}^{(r)}$ are the two distinct real roots of
$B_{k-1}x^{2}+(B_{k-2}-A_{k-1})x-A_{k-2}=0.$ (21)
Recalling that
$\alpha_{1}^{(r)}+\alpha_{2}^{(r)}=\frac{A_{k-1}-B_{k-2}}{B_{k-1}},$
we get that Condition (20) is equivalent to
$\frac{2A_{k-1}}{B_{k-1}}=\alpha_{1}^{(r)}+\alpha_{2}^{(r)}=\frac{A_{k-1}-B_{k-2}}{B_{k-1}},$
that can be rewritten as
$\frac{A_{k-1}+B_{k-2}}{B_{k+1}}=0.$
Therefore, (20) holds if and only if
$A_{k-1}+B_{k-2}=0.$ (22)
Notice that if the latter holds then $k\neq 1$, as otherwise it would imply
$a_{0}=A_{0}=-B_{-1}=0,$
that would correspond to the purely periodic continued fraction $\overline{0}]$. By hypothesis $\alpha_{1}^{(r)}$ and $\alpha_{2}^{(r)}$ are real and distinct, so that (21) has positive discriminant, i.e.
$D=(B_{k-2}-A_{k-1})^{2}+4A_{k-2}B_{k-1}>0.$
Rearranging the latter expression and exploiting (7) and (22), we obtain that ii)ii) is not satisfied if and only if
$D$ $=(B_{k-2}-A_{k-1})^{2}+4A_{k-2}B_{k-1}=$
$=(B_{k-2}+A_{k-1})^{2}-4A_{k-1}B_{k-2}+4A_{k-2}B_{k-1}=4\cdot(-1)^{k-1}>0.$
Hence, whenever $D>0$, (20) can hold only if the period $k$ is odd. The period length $k$ is always even for Browkin II, therefore condition ii)ii) of Theorem 15 is always satisfied for this algorithm. Moreover notice that, if the period length is odd in Algorithm (12), then all its partial quotients have negative valuations. The reason is that all the odd partial quotients in Algorithm (12) have negative valuation, therefore if $v_{p}(a_{i})=0$ for some even $i$, then by the periodicity also $a_{i+k}=a_{i}$ has zero valuation, that is not possible as $i+k$ is an odd index. Now, observe that $A_{k-1}=-B_{k-2}$ implies $v_{p}(A_{k-1})=v_{p}(B_{k-2})$. It is known that
$v_{p}(A_{k-1})$ $=v_{p}(a_{0})+v_{p}(a_{1})+\ldots+v_{p}(a_{k-1}),$
$v_{p}(B_{k-2})$ $=v_{p}(a_{1})+v_{p}(a_{2})+\ldots+v_{p}(a_{k-2}).$
For the latter equalities see, for example, [8, Lemma 1] and [36, Remark 4.1]. It follows that $v_{p}(A_{k-1})=v_{p}(B_{k-2})$ only if
$v_{p}(a_{0})+v_{p}(a_{k-1})=0.$ (23)
This equation can not be satisfied by Browkin I as, by construction, $v_{p}(a_{0})\leq 0$ and $v_{p}(a_{k-1})<0$, and we have seen that $k\neq 1$ if (22) holds. Therefore, (20) is never satisfied for this algorithm. It remains to prove that the hypothesis ii)ii) of Theorem 15 holds for Algorithm (12). We proved that Equation (20) implies that the period length $k$ is odd, in the case of $D>0$. However, we showed also that, for odd $k$, all partial quotients in Algorithm (12) have negative valuation. This implies that $v_{p}(a_{0})+v_{p}(a_{k-1})<0$, so that $A_{k-1}\neq-B_{k-2}$, as we have seen before for Browkin I. Therefore, (20) can never be satisfied for Browkin I, Browkin II and Algorithm (12). It follows that the condition ii)ii) of Theorem 15 is always satisfied for the $p$-adic continued fractions generated by these algorithms, and this proves the claim. ∎
From the computational results leading to Conjecture 13, we see that the pp–adic continued fraction of a quadratic irrational seems to always converge to a real number. Moreover, from Proposition 1, when its pp–adic continued fraction is periodic, it converges in $\mathbb{R}$ to a root of the same quadratic polynomial. However, in cases where periodicity is not recognized, the continued fraction seems to converge very neatly “somewhere else” in $\mathbb{R}$. This means that, for those pp–adic quadratic irrationals $\alpha^{(p)}$ that we believe to be non-periodic (based on brute-force search of the periodicity for a large number of steps), the continued fraction is approaching really well a real number that is completely different from $\alpha^{(r)}$ or $\overline{\alpha}^{(r)}$. Since this is a necessary condition for the periodicity of the pp–adic continued fraction of $\alpha^{(p)}$, by Proposition 1, proving the convergence to another real number would be an example of a non-periodic Browkin’s continued fraction. However, we are not able to exhibit such an example. Even if the continued fractions that are apparently non-periodic seem to converge really well to another real number, it remains a (very convincing, but heuristic) computational observation. All the computational results are listed in Section 6.
Also, we did not observe any continued fraction of $\alpha^{(p)}$ that seems to converge in $\mathbb{R}$ to $\alpha^{(r)}$ or $\overline{\alpha}^{(r)}$ and does not become periodic. If it does not show a period, the continued fraction is converging really well but really far from $\alpha$ in Euclidean absolute value. Therefore, it seems likely that Conjecture 2 is true, and solving it would provide a really nice characterization of periodic pp–adic continued fractions.
The difficult part of Conjecture 2 is to prove that if the $p$-adic continued fraction of the quadratic irrational $\alpha^{(p)}$ converges in $\mathbb{R}$ to either $\alpha^{(r)}$ or $\overline{\alpha}^{(r)}$, then it becomes periodic at some point. We leave the following two conjectures, that are supported by experimental computations (part of them is reported in Section 6).
Conjecture 16.
Let $a_{0},a_{1},\ldots]$, where $a_{n}\in\mathbb{Q}$ for all $n$. If $a_{0},a_{1},\ldots]$ converge simultaneously in $\mathbb{R}$ and $\mathbb{Q}_{p}$, respectively to a real quadratic irrational $\alpha^{(r)}$ and its image $\alpha^{(p)}$ inside $\mathbb{Q}_{p}$, then the continued fraction $a_{0},a_{1},\ldots]$ is periodic.
Conjecture 17.
Let $a_{0},a_{1},\ldots]$ the pp–adic continued fraction, obtained with either Browkin I, Browkin II or Algorithm (12), of a quadratic irrational $\alpha^{(p)}$ that has an image $\alpha$ inside $\mathbb{R}$. If $a_{0},a_{1},\ldots]$ converge to either $\alpha^{(r)}$ or $\overline{\alpha}^{(r)}$ in $\mathbb{R}$, then the continued fraction is periodic.
Conjecture 16 implies Conjecture 17, as we are giving more restrictive hypotheses. In fact, in Conjecture 17 we have a more precise shape of the rational numbers of the continued fractions, because they are “legal” partial quotients of Browkin I, Browkin II or Algorithm (12).
Example 18.
Let us consider $\sqrt{19}$ in $\mathbb{Q}_{5}$. The polynomial $x^{2}-19$ has a root, and hence two roots, in $\mathbb{Q}_{5}$. This is the first and simplest example of $\sqrt{D}$, with $D>0$ non-square integer that is a quadratic residue modulo $5$, such that Browkin I and Browkin II continued fractions seem to be non-periodic. The expansions of $\sqrt{14}$ and $\sqrt{21}$, that are the quadratic residues modulo $5$ nearest to $\sqrt{19}$, have expansions
$\sqrt{14}=$ $\left[2,-\frac{3}{5},\overline{-\frac{9}{5},-\frac{6}{5},\frac{166}{125},-\frac{6}{5},-\frac{9}{5},-\frac{8}{5}}\right],$
$\sqrt{21}=$ $\left[1,\frac{3}{5},\overline{\frac{3}{5},-\frac{4}{5},\frac{3}{5},-\frac{7}{5},\frac{26}{25},-\frac{7}{5}}\right],$
with Browkin I and
$\sqrt{14}=$ $\left[2,\overline{\frac{2}{5},-1,-\frac{1}{5},2,-\frac{1}{5},-1}\right],$
$\sqrt{21}=$ $\left[1,\overline{-\frac{2}{5},1,\frac{2}{5},2,\frac{1}{5},-2,\frac{2}{5},-2,\frac{1}{5},2,\frac{2}{5},1,-\frac{2}{5},2}\right],$
with Browkin II. For $\sqrt{19}$, we did not find any period with a search up to $50.000$ complete quotients. It seems really unlikely to have periodicity of small square roots for small primes $p$ with such a long period. In fact, “small” square roots tend to have short periods whenever they are periodic, as it happens for real continued fractions and as we can see for $\sqrt{14}$ and $\sqrt{21}$. This observation already suggests that is really unlikely for $\sqrt{19}$ to have a periodic continued fraction with Browkin I and Browkin II. Also, if we look at the convergents, we can notice the following. For the Browkin I expansion of $\sqrt{19}$,
$\frac{A_{10}}{B_{10}}$ $=1.35736553571026\ldots,$
$\frac{A_{100}}{B_{100}}$ $=1.35738766711068\ldots,$
$\frac{A_{1000}}{B_{1000}}$ $=1.35738766711068\ldots,$
$\frac{A_{5000}}{B_{5000}}$ $=1.35738766711068\ldots,$
$\frac{A_{10000}}{B_{10000}}$ $=1.35738766711068\ldots,$
and this approximation just gets better going forward with the expansion. It seems that the Browkin I continued fraction of $\sqrt{19}$ is converging to a real number which is, however, different from $\sqrt{19}\approx 4.36$. Proving that this expansion is not actually convergent to something near $4.36$ would be a counterexample for the periodicity of Browkin I, by Proposition 1. However, we did not manage to prove it for $\sqrt{19}$ or for any other quadratic irrational. Similar results are true for Browkin II, where the convergents of $\sqrt{19}$ are
$\frac{A_{10}}{B_{10}}$ $=2.57225268855495\ldots,$
$\frac{A_{100}}{B_{100}}$ $=1.89462102495469\ldots,$
$\frac{A_{1000}}{B_{1000}}$ $=1.89443989021177\ldots,$
$\frac{A_{5000}}{B_{5000}}$ $=1.89443989021177\ldots,$
$\frac{A_{10000}}{B_{10000}}$ $=1.89443989021177\ldots.$
Also Browkin II convergents of $\sqrt{19}$ seem clearly converging to some real number. This real number, also in this case, is different from $\sqrt{19}\approx 4.36$. With our algorithm (12), $\sqrt{19}$ has periodic expansion
$\sqrt{19}=\left[2,\overline{-\frac{2}{5},2,\frac{1}{5},-2,-\frac{2}{5},-\frac{12}{5},\frac{2}{5},-2,\frac{8}{25},2,\frac{1}{5},-1,-\frac{2}{5},-\frac{8}{5},\frac{2}{5},-2,\frac{12}{25},2,\frac{2}{5},-1}\right].$
For $\sqrt{39}$ we did not observed any period up to $50.000$ steps with Algorithm (12). As for Browkin’s algorithms, the convergents seem to approach very neatly some real number. In fact,
$\frac{A_{10}}{B_{10}}$ $=3.11589405199643\ldots,$
$\frac{A_{100}}{B_{100}}$ $=3.24461192422490\ldots,$
$\frac{A_{1000}}{B_{1000}}$ $=3.23880830293096\ldots,$
$\frac{A_{5000}}{B_{5000}}$ $=3.23880830293096\ldots,$
$\frac{A_{10000}}{B_{10000}}$ $=3.23880830293096\ldots,$
that is far from $\sqrt{39}\approx 6.24$, to which they should converge if the continued fraction became periodic.
It seems really likely, in Example 18, that the convergents of $\sqrt{19}$ and $\sqrt{39}$ are converging to something else rather than to the real value of $\sqrt{19}$ and $\sqrt{39}$. We can notice it also by looking at the difference between two consecutive partial quotients
$\left|\frac{A_{n+1}}{B_{n+1}}-\frac{A_{n}}{B_{n}}\right|=\frac{1}{\left|B_{n}\right|\left|B_{n+1}\right|}.$ (24)
The denominators of the convergents are generated as in (6), therefore
$B_{n+1}=a_{n+1}B_{n}+B_{n-1}.$
For the continued fractions provided by Browkin’s algorithms and Algorithm (12), the denominators $B_{n}$ are not always increasing. In fact, since the partial quotients are allowed to be negative, and they can be in general arbitrarily small, the sequence $B_{n}$ is not necessarily increasing. However, as it is true that $|B_{n}|$ can also be small, it does not usually happen.
In fact, in order to have a small partial quotient, for example equal to $\frac{1}{p^{5}}$ in Browkin I, at some step $n$, we should have
$v_{p}(\alpha_{n-1}-a_{n-1})=v_{p}(c_{1}p+c_{2}p^{2}+\ldots)=5.$
It means that $c_{1}=c_{2}=c_{3}=c_{4}=0$, where $c_{1},c_{2},c_{3},c_{4}\in\left\{-\frac{p-1}{2},\ldots,\frac{p-1}{2}\right\}$ are representatives modulo $p$. However, as explained in the next section, it is widely believed that the $p$-adic digits of irrational numbers are uniformly distributed in $\{0,\ldots,p-1\}$, and hence in $\left\{-\frac{p-1}{2},\ldots,\frac{p-1}{2}\right\}$. Therefore, the probability of having four consecutive 0’s is $\frac{1}{p^{4}}$. The dominating term of $B_{n}$ is $a_{1}\cdot a_{2}\cdot\ldots\cdot a_{n}$. Therefore, when $|a_{i}|>1$, we expect to have an exponential growth of the $B_{n}$. Also, especially for big values of $p$, it is very rare to have $|a_{i}|<1$. In particular, notice that if in the $p$-adic expansion of
$a_{i}=c_{-r}\frac{1}{p^{r}}+\ldots+c_{-1}\frac{1}{p}+c_{0},$
$|c_{0}|\geq 2$, then $|a_{i}|>1$.
This observation is the reason why, in the next section, we carry out a probabilistic argument.
5. A probabilistic approach to the real convergence
In this section, we use Assumption 3 of uniform distribution of the digits in the $p$-adic expansion of a quadratic irrational. The aim is to build a probabilistic argument for the non-periodicity of pp–adic continued fractions obtained with Browkin I, Browkin II and Algorithm (12). As already anticipated in the introduction, Assumption 3 is based on Borel’s normal number theorem [6].
Definition 19.
Let $\alpha\in\mathbb{R}$ and let $b\geq 2$ be an integer. Then $\alpha$ is normal in base $b$ if every digit of the expansion of $\alpha$ in base $b$ has density $\frac{1}{b}$. We say that $\alpha$ is normal, or absolutely normal, if it is normal in every base $b\geq 2$.
In [6], Borel proved the following result.
Theorem 20 (Borel’s Theorem).
Almost all real numbers are normal, i.e. the set of non-normal numbers has Lebesgue measure 0 in $\mathbb{R}$.
Borel’s Theorem is proved by applying Borel-Cantelli Lemma. It is known that the strategy of the proof can be applied equivalently to the pp–adic expansion of pp–adic numbers. Therefore, almost all pp–adic numbers are normal, i.e. every digit in $\mathbb{Z}/p\mathbb{Z}$ has density $\frac{1}{p}$. In particular, all irrational numbers, both in $\mathbb{R}$ and in $\mathbb{Q}_{p}$, are conjectured to be normal. However, apart from Borel’s existence result, only few numbers have been proved to be normal. For example, there is not a proof that $\sqrt{2}$, $\pi$ or $e$ are normal, although it is strongly considered true.
Remark 21.
By Assumption 3, it is implied also the uniform distribution of the coefficients of the partial quotients in both Browkin’s and Ruban’s continued fractions (see, e.g., the metric results in [37]).
In the following, for sake of simplicity, we focus on Browkin I algorithm. However, the same arguments can be adapted to obtain similar results also for Browkin II and Algorithm (12).
First of all, we start by computing the probability for a complete quotient of the Browkin I continued fraction to have a given pp–adic valuation.
Proposition 22.
Let $\alpha_{0}\in\mathbb{Q}_{p}$ be a quadratic irrational. Then
$\mathbb{P}(v_{p}(\alpha_{n})=-k)=\frac{p-1}{p^{k}},$ (25)
for all $n,k\geq 1$.
Proof.
The valuation of $\alpha_{n}$ is k-k if and only if
$\alpha_{n-1}-a_{n-1}=c_{k}p^{k}+\ldots.$
This is equivalent to ask that $c_{1}=c_{2}=\ldots=c_{k-1}=0$, that happens with probability
$\mathbb{P}(c_{1}=c_{2}=\ldots=c_{k-1}=0)=\frac{1}{p^{k-1}},$
and $c_{k}\neq 0$, that happens with probability
$\mathbb{P}(c_{k}\neq 0)=\frac{p-1}{p},$
and this proves the claim. ∎
Remark 23.
Notice that, in fact
$\sum\limits_{k=1}^{+\infty}\frac{p-1}{p^{k}}$ $=(p-1)\left(\sum\limits_{k=1}^{+\infty}\frac{1}{p^{k}}\right)=(p-1)\left(\sum\limits_{k=0}^{+\infty}\frac{1}{p^{k}}\right)-(p-1)=$
$=(p-1)\left(\frac{1}{1-\frac{1}{p}}\right)-(p-1)=p-(p-1)=1,$
so that (25) actually defines a distribution on the natural numbers.
Then, in the next proposition we compute the expected value of the valuation of a Browkin I complete quotient.
Proposition 24.
Let $\alpha_{0}\in\mathbb{Q}_{p}$ be a quadratic irrational. Then,
$\mathbb{E}(v_{p}(\alpha_{n}))=\frac{p}{p-1},$
for all $n\geq 1$.
Proof.
Let us notice that
$\mathbb{E}(v_{p}(\alpha_{n}))=\sum\limits_{k=1}^{+\infty}k\frac{p-1}{p^{k}},$
is a convergent series since
$\lim\limits_{k\rightarrow+\infty}\frac{(k+1)\frac{p-1}{p^{k+1}}}{k\frac{p-1}{p^{k}}}=\lim\limits_{k\rightarrow+\infty}\frac{k+1}{kp}=\frac{1}{p}<1.$
Then we compute
$\mathbb{E}(v_{p}(\alpha_{n}))$ $=\sum\limits_{k=1}^{+\infty}k\frac{p-1}{p^{k}}=(p-1)\sum\limits_{k=1}^{+\infty}\frac{k}{p^{k}}+(p-1)\sum\limits_{k=1}^{+\infty}\frac{1}{p^{k}}-(p-1)\sum\limits_{k=1}^{+\infty}\frac{1}{p^{k}}=$
$=(p-1)\sum\limits_{k=1}^{+\infty}\frac{k-1}{p^{k}}+(p-1)\sum\limits_{k=1}^{+\infty}\frac{1}{p^{k}}=$
$=(p-1)\sum\limits_{k=0}^{+\infty}\frac{k}{p^{k+1}}+(p-1)\left(\frac{1}{1-\frac{1}{p}}-1\right)=$
$=\frac{1}{p}\sum\limits_{k=0}^{+\infty}k\frac{p-1}{p^{k}}+1=\frac{1}{p}\sum\limits_{k=1}^{+\infty}k\frac{p-1}{p^{k}}+1.$
It means that, for $\mathbb{E}=\mathbb{E}(v_{p}(\alpha_{n}))$,
$\mathbb{E}=\frac{1}{p}\mathbb{E}+1,$
that is,
$\mathbb{E}=\mathbb{E}(v_{p}(\alpha_{n}))=\frac{p}{p-1},$
and the claim is proved. ∎
Remark 25.
Proposition 24 tells us that, especially for large prime $p$, having large negative valuation is extremely rare. This is what we expect, since large valuations are caused by several consecutive $0^{\prime}s$, and under Assumption 3 each zero coefficient appears with probability $\frac{1}{p}$.
Now we want to compute the expected size of Browkin I partial quotients inside $\mathbb{R}$, in order to predict the behaviour of the Euclidean absolute value of the convergents. Notice that the expected value of any partial quotient $a$ in Browkin I is $\mathbb{E}(a)=0$. This is due to the choice of the symmetric interval $\{-\frac{p-1}{2},\ldots,\frac{p-1}{2}\}$. Therefore, to have an estimate of the size of the partial quotients, we compute the expected value of its Euclidean absolute value.
Theorem 26.
Let $a$ be a Browkin I partial quotient and let $k\geq 1$. Then, the expected value of its Euclidean absolute value is
$\mathbb{E}(|a|\ |\ v_{p}(a)=-k)=\frac{p^{2(k+1)}-1}{4p^{2k+1}},$
and, in general,
$\mathbb{E}(|a|)=\frac{p}{4}\left(1-\frac{1}{p^{2}(p^{2}+p+1)}\right).$
Proof.
The expected value of the constant term $c_{0}$ is 0, but in absolute value it is
$\mathbb{E}(|c_{0}|)$ $=\frac{1}{p}(0+2+\ldots+p-1)=\frac{2}{p}\left(1+\ldots+\frac{p-1}{2}\right)=$
$=\frac{2}{p}\left(\frac{p-1}{2}\right)\left(\frac{p+1}{2}\right)\frac{1}{2}=\frac{p^{2}-1}{4p}.$
The expected value of $|c_{0}+c_{1}\frac{1}{p}|$ is composed by the case where $c_{0}=0$, i.e.
$\frac{2}{p}\left(1+\ldots+\frac{p-1}{2}\right),$
and the case where $c_{0}\neq 0$. When $c_{0}\neq 0$, the positive elements $c_{0}+c_{1}\frac{1}{p}$ are exactly the ones with positive $c_{0}\in\left\{0,\ldots,\frac{p-1}{2}\right\}$. For $c_{0}\neq 0$, all the elements $c_{0}+c_{1}p$ and $c_{0}-c_{1}p$ cancel out, and all of them are counted twice (the positive and the negative one). Therefore, for all choices of $c_{0}\neq 0$ we have a summand of the kind $2pc_{0}$, since the tails with $c_{1}$ eliminate and there are exactly $p$ summands. Now we can compute
$\mathbb{E}\left(\left|c_{0}+\frac{c_{1}}{p}\right|\right)$ $=\frac{1}{p^{2}}\left(\frac{2}{p}\left(1+\ldots+\frac{p-1}{2}\right)+2p\left(1+\ldots+\frac{p-1}{2}\right)\right)=$
$=\frac{1}{p^{2}}\cdot\frac{p-1}{2}\cdot\frac{p+1}{2}\cdot\frac{1}{2}\left(\frac{2}{p}+2p\right)=$
$=\frac{p^{2}-1}{8p^{2}}\cdot\frac{2+2p^{2}}{p}=\frac{p^{4}-1}{4p^{3}}.$
In the general case, with a similar reasoning we obtain
$\mathbb{E}\left(\left|c_{0}+\frac{c_{1}}{p}+\ldots+\frac{c_{k}}{p^{k}}\right|\right)$ $=\frac{1}{p^{2}}\mathbb{E}\left(\left|c_{1}+\ldots+\frac{c_{k}}{p^{k-1}}\right|\right)+\frac{2p^{k}}{p^{k+1}}\left(1+\ldots+\frac{p-1}{2}\right)=$
$=\frac{1}{p^{2}}\mathbb{E}\left(\left|c_{1}+\ldots+\frac{c_{k}}{p^{k-1}}\right|\right)+\frac{2p^{k}(p^{2}-1)}{8p^{k+1}}=$
$=\frac{1}{p^{2}}\mathbb{E}\left(\left|c_{1}+\ldots+\frac{c_{k}}{p^{k-1}}\right|\right)+\frac{p^{2}-1}{4p}.$
The reason why $\frac{1}{p^{2}}$ appears in the first summand, is that, when $c_{0}=0$, we can collect $\frac{1}{p}$ from the sum and, moreover, the expected value on $k+1$ terms is divided by $p$ one time more than the expected value on $k$ terms. For the second summand, as before we have two times every value, every $k$-tuple cancel and they are exactly $p^{k}$, so we multiply that for every value of $c_{0}$ in $\{0,\ldots,\frac{p-1}{2}\}$. Now, if we call
$\mathbb{E}_{k}=\mathbb{E}\left(\left|c_{0}+\frac{c_{1}}{p}+\ldots+\frac{c_{k}}{p^{k}}\right|\right),$
we have:
$\mathbb{E}_{k}$ $=\frac{p^{2}-1}{4p}+\frac{1}{p^{2}}\mathbb{E}_{k-1}=\frac{p^{2}-1}{4p}+\frac{p^{2}-1}{4p}\cdot\frac{1}{p^{2}}+\frac{1}{p^{4}}\mathbb{E}_{k-2}=$
$=\frac{p^{2}-1}{4p}\left(1+\frac{1}{p^{2}}+\frac{1}{p^{4}}+\ldots+\frac{1}{p^{2(k-1)}}\right)+\frac{1}{p^{2k}}\mathbb{E}_{0}=$
$=\frac{p^{2}-1}{4p}\left(1+\frac{1}{p^{2}}+\frac{1}{p^{4}}+\ldots+\frac{1}{p^{2(k-1)}}+\frac{1}{p^{2k}}\right)=$
$=\frac{p^{2}-1}{4p}\left(\frac{p^{2(k+1)}-1}{p^{2k}(p^{2}-1)}\right)=\frac{p^{2(k+1)}-1}{4p^{2k+1}}.$
Then, we can compute
$\mathbb{E}(|a|)$ $=\sum\limits_{k=1}^{+\infty}\mathbb{E}(|a|\ |v_{p}(a)=-k)\mathbb{P}(v_{p}(a)=-k)=\sum\limits_{k=1}^{+\infty}\frac{p^{2(k+1)}-1}{4p^{2k+1}}\cdot\frac{p-1}{p^{k}}=$
$=\frac{p-1}{4}\sum\limits_{k=1}^{+\infty}\frac{p^{2k+2}-1}{p^{3k+1}}=\frac{p-1}{4}\left(\sum\limits_{k=1}^{+\infty}\frac{1}{p^{k-1}}-\sum\limits_{k=1}^{+\infty}\frac{1}{p^{3k+1}}\right)=$
$=\frac{p-1}{4}\left(\sum\limits_{k=0}^{+\infty}\frac{1}{p^{k}}-\frac{1}{p^{4}}\sum\limits_{k=0}^{+\infty}\frac{1}{p^{3k}}\right)=\frac{p-1}{4}\left(\frac{1}{1-\frac{1}{p}}-\frac{1}{p^{4}}\cdot\frac{1}{1-\frac{1}{p^{3}}}\right)=$
$=\frac{p-1}{4}\left(\frac{p}{p-1}-\frac{1}{p}\cdot\frac{1}{p^{3}-1}\right)=\frac{1}{4}\left(p-\frac{1}{p}\cdot\frac{1}{p^{2}+p+1}\right)=$
$=\frac{p}{4}\left(1-\frac{1}{p^{2}(p^{2}+p+1)}\right),$
and this concludes the proof. ∎
Remark 27.
Let us notice that we always expect $|a|>1$ for all primes $p$ except $p=3$. Moreover, the expected size of every partial quotients grows linearly with $p$:
$p=13\ \ \ \$$\Longrightarrow\ \ \ \ \mathbb{E}(|a|)\approx 3.25,$
$p=43\ \ \ \$$\Longrightarrow\ \ \ \ \mathbb{E}(|a|)\approx 10.75,$
$p=211\ \ \ \$$\Longrightarrow\ \ \ \ \mathbb{E}(|a|)\approx 52.75,$
$p=839\ \ \ \$$\Longrightarrow\ \ \ \ \mathbb{E}(|a|)\approx 209.75.$
Example 28.
For $p=7823$, the convergents of the expansion of $\sqrt{15648}$ with Browkin I are:
$\frac{A_{10}}{B_{10}}$ $=3339.99956244164\ldots,$
$\frac{A_{100}}{B_{100}}$ $=3339.99956244164\ldots,$
$\frac{A_{1000}}{B_{1000}}$ $=3339.99956244164\ldots,$
$\frac{A_{5000}}{B_{5000}}$ $=3339.99956244164\ldots,$
$\frac{A_{10000}}{B_{10000}}$ $=3339.99956244164\ldots.$
Indeed, the continued fraction seems to rapidly converge to a real limit, which is different from $\sqrt{15648}\approx 125.09$. In this case, it is really unlikely for the convergents to deviate from that value, due to Equation (24). In fact, we expect $|B_{n}|$ to grow exponentially as
$|B_{n}|\sim|a_{1}|\ldots|a_{n}|,$
where the expected value of each partial quotient is, by Theorem 26,
$\mathbb{E}(|a|)\approx 1955.75.$
Finally, based on these results and the computational analysis performed in the next section we leave Conjecture 2. In practice, by Conjecture 2, even if cancellation can happen and the convergents can be arbitrarily oscillating, this is really unlikely, as the sequence $|B_{n}|$ tends to have an exponential growth.
6. Computations
In the following tables, we list the approximations of $\sqrt{D}$, having both a $p$-adic and a real image, provided by the $1000$th convergent of Browkin I, Browkin II and Algorithm (12) pp–adic continued fractions in $\mathbb{Q}_{5}$ and $\mathbb{Q}_{43}$, distinguishing the periodic and the apparently non-periodic cases. It is possible to notice that, when the pp–adic continued fraction of $\sqrt{D}$ is periodic, the convergents approach very well and very soon the real limit, that is $\sqrt{D}$ by Proposition 1. In the cases where periodicity is not detected, the continued fraction seems to approach a real limit different from $\sqrt{D}$. This is a behaviour that we have observed in all the examples of our analysis. The SageMath code used for the computational part is publicly available [^1] https://github.com/giulianoromeont/p-adic-continued-fractions and contains the implementations of the main algorithms for pp–adic continued fractions.
Table 1. Non-periodic Browkin I expansions within $1000$ steps in $\mathbb{Q}_{5}$.
D $(A_{1000}/B_{1000})^{2}$
19 1.84226
26 0.42758
29 1.56850
31 3.85808
39 1.46700
41 3.17409
44 5.17107
46 0.23921
51 0.04092
56 0.00378
59 6.02309
61 1.76172
66 6.74804
71 1.83277
79 6.64195
84 2.60241
86 10.4167
89 15.3162
91 5.00953
96 0.16662
101 1.78169
106 2.94191
109 7.64688
114 14.9227
116 0.10316
124 0.87815
126 1.38815
129 1.67262
131 0.55175
134 7.69174
136 0.63808
139 1.46264
141 6.41153
146 2.86117
149 2.85846
151 0.00054
Table 2. Periodic Browkin I expansions within $1000$ steps in $\mathbb{Q}_{5}$.
D $(A_{1000}/B_{1000})^{2}$
6 6.00005
11 11.00049
14 14.00031
21 20.99472
24 24.00020
34 34.00056
54 54.00045
69 68.99960
74 73.99268
76 75.99829
94 93.99884
99 99.00051
104 104.00124
111 110.99886
119 119.00191
Table 3. Non-periodic Browkin II expansions within $1000$ steps in $\mathbb{Q}_{5}$.
D $(A_{1000}/B_{1000})^{2}$
19 3.58875
39 11.61105
41 0.24482
44 1.84987
46 0.18939
59 0.08105
66 3.49091
71 0.00555
74 5.30380
76 2.88456
86 0.13089
89 29.0251
94 13.4146
96 7.01349
99 0.03798
101 2.46992
106 0.00040
119 4.90356
124 18.46678
131 3.90023
134 7.01402
141 3.03630
146 0.00366
151 5.06295
154 18.45303
159 14.69725
164 0.91049
166 0.42797
174 6.30210
179 0.06295
181 0.52794
184 2.80529
186 14.3330
189 2.5363
191 26.1437
194 2.6614
199 17.37722
Table 4. Periodic Browkin II expansions within $1000$ steps in $\mathbb{Q}_{5}$.
D $(A_{1000}/B_{1000})^{2}$
6 5.99956
11 10.99983
14 13.99957
21 21.92018
24 24.00020
26 25.99980
29 28.99930
31 30.99928
34 34.00056
51 50.99959
54 53.99898
56 55.99977
61 61.00078
69 38.92137
79 78.99832
84 84.75043
91 91.00015
104 104.0012
109 72.80696
111 110.9525
114 113.9983
116 115.9993
126 125.9983
129 128.9996
136 136.0022
139 138.9993
149 148.9888
156 155.9976
161 160.8153
171 170.9974
176 175.9973
Table 5. Non-periodic Algorithm (12) expansions within $1000$ steps in $\mathbb{Q}_{5}$.
D $(A_{1000}/B_{1000})^{2}$
39 10.4898
41 0.17799
46 0.12453
66 0.07997
69 2.10163
71 0.00543
79 1.61061
86 0.13053
89 13.0321
101 2.4583
106 4968.642
109 85110.827
114 26.4699
124 18.35951
131 3.6963
136 8.6183
141 4.7646
146 1.04080
149 0.02709
151 2.85170
154 19.7118
159 16.6455
161 10.9441
164 0.7972
171 0.0039
174 19.1686
179 0.1908
181 31449.44
184 2.80562
186 22.7023
189 15.1905
194 2.6604
199 15.142
Table 6. Periodic Algorithm (12) expansions within $1000$ steps in $\mathbb{Q}_{5}$.
D $(A_{1000}/B_{1000})^{2}$
6 5.9995
11 10.9998
14 13.9995
19 19.000
21 21.9201
24 24.0002
26 25.999
29 28.9993
31 30.9992
34 34.0005
44 43.9900
51 50.9995
54 54.0004
56 55.9997
59 59.0008
61 61.0007
74 73.9995
76 75.9982
84 84.7504
91 90.9982
94 93.9988
96 95.9988
99 99.0005
104 104.001
111 110.996
116 115.999
119 117.887
126 125.998
129 129.001
134 133.999
139 138.999
156 155.997
166 142.515
176 174.195
191 190.997
Table 7. Non-periodic Browkin I expansions within $1000$ steps in $\mathbb{Q}_{43}$.
D $(A_{1000}/B_{1000})^{2}$
6 44.7453
10 222.0129
11 436.6093
13 402.1308
14 98.7578
15 146.7199
17 367.3585
21 62.1133
23 326.9225
24 197.4165
31 296.631
38 82.5499
40 166.265
41 254.491
44 0.9057
47 4.3664
52 8.1533
53 227.689
54 438.270
56 403.029
57 98.7161
58 146.785
59 15.5031
60 356.042
66 330.428
67 192.820
68 26.185
74 286.845
79 36.648
83 174.448
84 247.882
87 0.907
90 4.373
92 43.991
95 8.1618
96 222.173
97 445.349
99 424.615
Table 8. Periodic Browkin I expansions within $1000$ steps in $\mathbb{Q}_{43}$.
D $(A_{1000}/B_{1000})^{2}$
35 35.000
78 78.000
187 186.997
Table 9. Non-periodic Browkin II expansions within $1000$ steps in $\mathbb{Q}_{43}$.
D $(A_{1000}/B_{1000})^{2}$
10 115.362
11 672.676
13 331.167
15 57.450
23 1301.16
31 220.368
35 60.478
40 240.74
53 300.467
54 358.488
56 569.003
58 84.0687
60 464.238
66 188.872
67 411.809
74 352.617
83 106.822
96 135.124
97 299.172
99 318.033
101 210.934
103 676.431
109 445.940
110 133.830
117 389.825
126 224.244
127 325.751
130 9.8125
133 4.7093
139 162.889
142 492.067
145 34.729
146 1808.22
153 309.851
154 2.678
160 1137.27
Table 10. Periodic Browkin II expansions within $1000$ steps in $\mathbb{Q}_{43}$.
D $(A_{1000}/B_{1000})^{2}$
6 5.9995
14 13.999
17 17.006
21 20.999
24 23.999
38 37.999
41 40.999
44 43.999
47 46.999
52 51.999
57 56.999
59 58.999
68 67.999
78 77.998
79 78.998
84 83.999
87 86.998
90 89.999
92 91.998
95 94.998
102 101.999
107 106.998
111 110.998
122 121.998
124 123.999
135 134.998
138 137.999
140 139.998
143 142.998
150 149.998
152 151.999
Table 11. Non-periodic Algorithm (12) expansions within $1000$ steps in $\mathbb{Q}_{43}$.
D $(A_{1000}/B_{1000})^{2}$
10 101.090
11 675.849
13 331.167
15 57.4503
23 1301.16
31 220.368
35 60.478
40 240.740
53 300.467
54 358.488
56 569.003
58 84.0687
60 464.238
66 188.872
67 411.408
74 352.617
83 106.822
96 135.124
97 299.172
99 318.033
101 210.934
103 676.431
109 445.940
110 133.839
117 389.821
126 224.244
127 325.751
130 9.8125
133 4.7093
139 162.889
142 216.787
145 34.7298
146 1808.222
153 309.851
154 2.6791
160 1137.254
Table 12. Periodic Algorithm (12) expansions within $1000$ steps in $\mathbb{Q}_{43}$.
D $(A_{1000}/B_{1000})^{2}$
6 5.9995
14 13.9995
17 17.0065
21 20.9993
24 23.9992
38 37.9998
41 40.9996
44 43.9993
47 46.9992
52 51.9999
57 56.9994
59 58.9992
68 67.9998
78 77.9989
79 78.9983
84 83.9990
87 86.9985
90 89.9993
92 91.9987
95 94.9981
102 101.999
107 106.998
111 110.998
122 121.998
124 123.999
135 134.998
138 137.999
140 139.998
143 142.998
150 149.998
152 151.999
Acknowledgments
The author is a member of GNSAGA of INdAM.